OpenStudy (anonymous):

What is the outcome of 16+18v+v^2 when factored completely?

OpenStudy (jdoe0001):

\(\bf 16+18v+v^2\implies v^2+16+18v\) I gather I'd use the quadratic formula

OpenStudy (anonymous):

I did a few other ones like this and got them but like idk this one I can’t figure to at all no matter how I try to do it :/

OpenStudy (jdoe0001):

have you covered the quadratic formula yet? because, the factoring doesn't give plain integers, I'd think it gives rationals so in that case splitting them by means of the constant factors, may not be as simple

OpenStudy (jdoe0001):

so if yo haven't covered the quadratic formula how are you supposed to get the factors, or how are you meant to get them? what section of the chapter is that covering that is

OpenStudy (anonymous):

\frac{ x=-b (+/-)sqrt{b ^{2}-4ac} }{2a } You probably havent gotten it yet because it doesnt factor that nicely. There are quite a few ways to factor a quadratic equation, but the quadratic formula is the only sure fire way to get the solutions youre looking for, as long as you do the arithmetic correctly. Honestly, it looks scarier than it is. It's more time consuming then anything. So, once you have the equation in standard form:\[v ^{2}+18v+16\] You can 'assign'them their letters. A=1 (V is the leading coefficient, and the coefficient is one) B=18 C=16 Then, using these numbers, you plug them in :D \[\frac{ x=-18(+/-)\sqrt{18 ^{2}-4(1)(16)} }{2(1) }\] That's pretty much it. You'll get two solutions because of the "plus or minus" thats in there. If you need any more help i would be happy to work through it some more with you